OpenStudy (anonymous):

If an angle theta increases uniformly, find the smallest positive value of theta for which tan theta increases 8 times as fast as sin theta

OpenStudy (anonymous):

Would it be something like this? \[\Large \frac{d}{d\theta}tan(\theta) =8\cdot \frac{d}{d\theta}sin(\theta) \] Where \[\Large \frac{d}{d\theta}tan(\theta) > 0\] What is the topic of the curriculum?

OpenStudy (anonymous):

Diffrentiation with respect to time.. @wio

OpenStudy (anonymous):

So do you think you can find those derivatives, and then solve for \(\theta\)?

OpenStudy (anonymous):

\[\sec^2 \theta =8 \cos \theta]

OpenStudy (anonymous):

\[\sec^2 \theta =8 \cos \theta\]

OpenStudy (anonymous):

Now, what is \(\sec^2(\theta )\) in terms of \(\sin(\theta) \) and \(\cos(\theta)\)?

OpenStudy (anonymous):

i know is sec = 1/ cos

OpenStudy (anonymous):

Do you still need help solving for \(\theta\)?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

Ok so we have \[\Large \frac{1}{\cos^2(\theta)} = 8\cdot \cos(\theta)\]How can we isolate \(\theta \) further?

OpenStudy (anonymous):

what will happen next?? no idea. -_-

OpenStudy (anonymous):

How about we multiply both sides by \(\cos^2(\theta)\)? Try that.

OpenStudy (anonymous):

then it will become 1= 8 cos^3 theta ??

OpenStudy (anonymous):

Yes! So what about getting rid of the coefficient?

OpenStudy (anonymous):

1/8 = cos ^3 theta ??

OpenStudy (anonymous):

Now it's just algebra. We learned that long ago.

OpenStudy (anonymous):

How do you get rid of an exponent?

OpenStudy (anonymous):

hmm i dont know can u help about it?

OpenStudy (anonymous):

Why don't you take the cubed root of both sides?

OpenStudy (anonymous):

oww okay I get it :) THanks

OpenStudy (anonymous):

Just remember that you want the smallest positive \(\theta \), and that \(\cos(\theta)\) must also be positive since they should be increasing.

OpenStudy (anonymous):

Otherwise there would be many solutions!

OpenStudy (anonymous):

I get theta = 60 is that correct?

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Thank you :)

hartnn (hartnn):

Great explanation @wio :)